Find the equation of the line that has a slope of -4 and passes through the point (2,3)

Answer by checkley71(8403)

Y=mX+b
Y=-4X+b REPLACING X & Y FOR 2 & 3 WE GET
3=-4*2+b
3=-8+b
3+8=b
11=b THE Y INTERCEPT
Y=-4X+11 IS THE EQUATION FOR THIS LINE